OpenStudy (donnie1999):

-------- WILL MEDAL -------- -------- HELP PLEASE -------- Find the x-intercepts for the parabola defined by the equation below. y = x^2 + x - 12

OpenStudy (donnie1999):

@Nixy i got (-4,0) and (3,0)

OpenStudy (butterflydreamer):

x-intercepts are when y = 0 x^2 + x - 12 = 0 *factorise* (x - 3) ( x + 4) = 0 x = 3, -4 Therefore x-intercepts are at: (3, 0) and ( -4, 0) Sooo, you're correct :) excellent work :D!